OpenStudy (anonymous):

write in scientific notation and std form (5 x 10^2) (7 x 10^-4)

OpenStudy (anonymous):

Std = 0.35 SN = 3.5x10^-1

OpenStudy (anonymous):

How did you get to the -1 exponent? That is where I got lost

OpenStudy (anonymous):

\[5\times 7=35\] \[10^{2}\times 10^{-4}=10^{-2}\]so you have \[35\times 10^{-2}\] but this is not scientific notation, because 35 is larger than ten

OpenStudy (anonymous):

so you have to change 35 to 3.5

OpenStudy (anonymous):

since you moved decimal one place to the left, you have to compensate in the exponent. \[35\times 10^{-2}=35\times 0.01=0.35\] to put this number in scientific notation you need \[3.5\] so you write \[3.5\times 10^{-1}\] which is \[3.5\times 0.1=0.35\]